# Odd-parity effect and scale-dependent viscosity in atomic quantum gases

### v1

@@ Abstract
It has recently been predicted that two-dimensional electron gases possess an anomalous “tomographic” transport regime outside of the traditional collisionless and hydrodynamic limits, but an experimental confirmation has been elusive so far. This anomalous regime is marked by the appearance of an odd-even effect in the quasiparticle lifetimes where deformations of the Fermi surface with odd-parity become long-lived in comparison to even-parity ones. In this work, we establish neutral atomic quantum gases as an alternative platform to reveal this new transport regime and demonstrate an odd-even effect in the normal phase of two-component Fermi gases. By diagonalizing the Fermi liquid collision integral, we identify odd-parity modes with anomalously long lifetimes below temperatures $T\leq 0.1T_{F}$, which is within the reach of current cold atom experiments. In a marked difference from condensed matter setups, we show that the odd-even effect in neutral gases is widely tunable with interactions along the BCS-BEC crossover and suppressed on the BEC side where the Fermi surface is destroyed. We propose the damping rate of quadrupole oscillations as an experimental signature of the long-lived odd-parity modes. The damping rate is set by the shear viscosity, which for finite trap confinement is dominated by odd-parity modes and thus anomalous enhanced compared to the hydrodynamic limit. Furthermore, a full computation of the shear viscosity within Fermi liquid theory shows that the magnitude of the odd-even effect depends on the particle number and is particularly pronounced in mesoscopic Fermi gases. Our findings suggest that the hydrodynamic behavior of neutral degenerate quantum gases is much richer than previously thought and should include additional long-lived modes.
@@ Introduction
Thermodynamic and transport properties of the normal state of metals as well as 3He are well described by a non-interacting Fermi gas model [1, 2]. The microscopic justification of this surprisingly simple description, despite nominally strong interactions, is provided by Landau’s theory of the Fermi liquid [3, 4], which describes thermodynamic and transport properties in terms of a collection of weakly interacting quasiparticles [5, 6, 7]. The validity of this picture relies on the presence of a Fermi surface and concomitant Pauli blocking, which strongly restricts the available phase space for quasiparticle scattering, leading to well-defined quasiparticles with lifetimes that diverge quadratically at low temperatures, $\tau\sim 1/T^{2}$. In recent years, Fermi liquid theory has also been applied to a variety of ultracold quantum gas setups, such as minority polarons in strongly spin-imbalanced Fermi gases [8, 9, 10, 11], the normal phase of balanced Fermi gases above the superfluid transition [12, 13, 14, 15], transport and collective modes [16] and even few-body systems [17]. For transport measurements, quasiparticle interactions lead to internal dissipative processes with relaxation times that set the magnitude of transport coefficients like the shear viscosity. The general expectation is that these transport times are equal to the quasiparticle lifetimes.
@@
Figure 1: (a) Nonequilibrium momentum-space distribution for different temperatures and (b) relaxation rates $\beta\hbar/\tau_{m}$ of the first collective even-parity Fermi surface deformation with angular momentum $m=2$ (blue) and the first odd-parity deformation with $m=3$ (orange). Even (odd) parity refers to an equal (opposite) deformation of the distribution at opposite momenta. With decreasing temperature and increasing density (i.e., increasing $\beta\mu$), the equilibrium Fermi surface becomes sharper and Pauli blocking diminishes the available phase space for scattering, increasing the lifetime of the mode. The conventional Fermi liquid relaxation rate $\sim T^{2}/\mu$ of the even-parity mode is set by an angular redistribution on the Fermi surface of pairs with equal and opposite momenta. Such a process does not relax the odd-parity modes, however, which have a significantly slower relaxation rate $\sim T^{4}/\mu^{3}$. For interacting Fermi gases along the BCS-BEC crossover, the discrepancy of odd and even-parity lifetimes is present on the BCS side of the crossover, for which the chemical potential is positive, $\beta\mu>0$ (the grey dashed line indicates the non-interacting chemical potential at a moderate small temperature $T/T_{F}=0.1$ for reference). The effect is then reduced with increasing pairing interaction and vanishes on the BEC side where $\beta\mu<0$. (c) Real-space dynamics and (d) damping rate $\Gamma_{Q}$ as a function of trap frequency $\omega_{r}$ of the quadrupole mode in a harmonic trap. The quadrupole mode oscillates with frequency $\omega_{Q}=\sqrt{2}\omega_{r}$ and involves a shear flow in real space. The real space density profile is shown in the insets and sets the equilibrium size $R$. The damping rate of the quadrupole mode is controlled by the shear viscosity spectral function $\eta(\omega,q)$, which is proportional to the relaxation time of a collective shear deformation of the Fermi surface in momentum space. In a bulk system (shallow trap), this deformation corresponds to the hydrodynamic even-parity mode with $m=2$, but at finite confinement odd-parity modes contribute, and the increased odd-parity lifetime leads to an increased damping rate (orange line) compared to the hydrodynamic prediction (blue line). This damping provides evidence of the enhancement of the Fermi liquid lifetime.
@@
Recent theoretical work in the context of interaction-dominated electron gases in two dimensions has provided evidence to the contrary and indicates the existence of deformations of the Fermi surface with significantly longer lifetimes [18, 19, 20, 21, 22]: Indeed, a standard Fermi liquid scaling $\tau_{\rm e}\sim 1/T^{2}$ is argued to only apply to a subclass of collective quasiparticle deformations—dubbed even-parity modes—while the remaining modes—dubbed odd-parity modes—decay significantly more slowly as $\tau_{\rm o}\sim 1/T^{4}$. This leads to the exciting perspective that, at low temperatures, these long-lived modes impact the hydrodynamic description of the gas (in addition to modes that are linked to conserved currents and have infinite lifetime), such that the transport properties of Fermi liquids become richer than commonly assumed. For this reason, there is now intense effort to identify experimental signatures of odd-parity transport, or “tomographic transport”, in electron Fermi liquids [19, 23, 24, 25, 26, 27]. Quite generally, an isolated odd-parity response that is not dominated by even-parity modes appears in transverse probes at finite wavelengths, such as the shear viscosity [19] or transverse collective modes [23]. However, it has so far not been possible to detect such transverse probes in Fermi liquids of electrons, while in 3He, measurements of transverse sound have been inconclusive [28, 29]. Quantum gases offer an entirely separate way of inducing transverse dynamics by manipulating the real-space confinement of an external trap. Moreover, on a microscopic level the odd-even effect is linked to the Fermi statistics and Pauli blocking, and one should thus expect (and as we confirm theoretically) that this effect is not only present for electron gases but also in charge-neutral Fermi liquids. The purpose of this work is thus to initiate a search for the odd-even effect in neutral quantum gases.
@@
The key results of our work, illustrated in Fig. 1, are twofold: First, we establish that an identical enhancement as for electron gases in the lifetime of odd-parity deformations of the Fermi surface [Fig. 1(a)] also exists for neutral atomic Fermi quantum gases. Using an exact diagonalization of the linearized kinetic collision integral (see methods), we find that these lifetimes differ significantly within temperature ranges accessible in current cold atom experiments [Fig. 1(b)]. Moreover, in a marked difference from electron gases, we demonstrate that the effect is widely tunable with interaction strength along the BCS-BEC crossover and vanishes near resonance and on the BEC side of the crossover where the chemical potential is negative and the Fermi surface disappears. Second, we propose as an experimental signature of the odd-even effect the damping rate of collective quadrupole oscillations in a harmonic trap, which involves a shear flow of atoms in real space [Fig. 1(c)]. The damping rate of the oscillation is set by the shear viscosity, which is evaluated at a finite wavenumber set by the trap curvature. As discussed above, this implies that the damping rate is proportional to the relaxation time of the momentum space deformation that is induced by the shear flow in the harmonic trap. For a shallow trap, the standard hydrodynamic result for the shear viscosity involves an even-parity deformation with conventional Fermi liquid lifetime, while for strong confinement, odd-parity modes dominate. An increased damping is thus a signature of the enhanced odd-parity lifetime, as the shear viscosity is proportional to the lifetime. Indeed, a full evaluation of the momentum- and frequency-dependent shear viscosity spectral function [Fig. 1(d)] shows for intermediate trap frequencies $\omega_{r}$ a strong increase of the damping rate compared to the hydrodynamic prediction for realistic system parameters. Moreover, as will be discussed below, the effect is even further enhanced both at low temperatures and for small atom numbers, i.e., for mesoscopic Fermi gases. Overall, these findings imply that in experiments with cold gases, the odd parity effect is widely tuned in a characteristic manner with interaction strength, temperature, harmonic trap frequency, and particle number. Ultracold Fermi gases are thus an ideal platform to study the anomalous relaxation dynamics of collective modes, and it should be readily possible to establish the odd-even effect in the quasiparticle lifetimes by comparing the damping of the quadrupole mode at different experimental parameter values.
@@ Results
Quasiparticle lifetimes We first establish the odd-even effect in the lifetimes of odd and even parity modes at low temperatures on the BCS side of the normal region of the BCS-BEC crossover (where a well-defined Fermi surface exists) and show that it vanishes along the crossover. Intuitively, this is linked to the fact that fermions along the crossover form increasingly tightly bound dimers, which reduces the chemical potential and destroys the Fermi surface, and with it the Pauli blocking that is at the heart of the odd-even effect. For 2D atomic gases, interactions are short-ranged and described by a delta-function with strength $\hbar^{2}\tilde{g}/m^{*}$ with a dimensionless parameter $\tilde{g}=-2\pi/\ln(k_{F}a_{2})$, where $k_{F}$ is the Fermi wavevector, $m^{*}$ is the mass of the atoms and $a_{2}$ is the 2D scattering length. At a given temperature and density, the interactions interpolate between the BCS side [$\tilde{g}<0$ or $\ln(k_{F}a_{2})>0$] and the BEC side [$\tilde{g}>0$ or $\ln(k_{F}a_{2})<0$], tuned by either Feshbach  [30] or confinement-induced resonances [31].
Figure 2: (a) Shear viscosity spectral function in units of $\hbar n$ as a function of wave number $q$ and frequency $\omega$ (both in thermal units with $\lambda_{T}=\sqrt{2\pi\beta\hbar^{2}/m^{*}}$). The white solid (dashed) line describes the path in the parameter space ($q=1/R,\omega=\omega_{Q}$) explored by tuning the trap confinement for $N=10^{3}$ ($N=10^{6}$) atoms. (b)-(d): Ratio of the shear viscosity spectral function $\eta(\omega=\omega_{Q},q=1/R)$ and the hydrodynamic shear viscosity $\eta_{HD}=\eta(\omega_{Q},0)$ as a function of (b) trap frequency, (c) interaction strength, and (d) particle number. Unless otherwise specified, we use an atom number $N=10^{3}$ and interaction strength $\ln(k_{F}a_{2})=0.63$. In each case, we observe a pronounced deviation from the hydrodynamic prediction for smaller atom number, interactions that are deeper in the BCS limit, and for weaker harmonic traps (or lower temperatures), respectively.
The fundamental quantity to describe are the relaxation rates of collective deviations in the quasiparticle distribution from local equilibrium (which is given by a Fermi-Dirac distribution $n_{\rm FD}({\bf p})$). These relaxation rates are defined as the eigenvalues of the linearized collision integral (see methods),
@@
Due to rotational invariance of the system, the eigenmodes $\psi_{m}$ carry a definite angular momentum quantum number $m=0,\pm 1,\pm 2...$. Modes with even (odd) $m$ have even (odd) parity. An additional radial quantum number (not indexed explicitly here) describes the number of radial nodes in the eigenfunction, with faster relaxation (i.e., larger $1/\tau_{m}$) for modes with increasing node number [21]. Due to conservation of particle number, energy, and momentum, the lowest eigenvalues with $m=0$ and $m=\pm 1$ are zero. The leading finite even-parity mode has then angular momentum $m=\pm 2$ while the leading odd-parity mode has angular momentum $m=\pm 3$ [21]. The corresponding momentum space deformations are shown in Fig. 1(a) for three different temperatures. The $m=2$ mode sets the dominant contribution to the hydrodynamic shear viscosity [32, 33, 34], while the second mode will give the leading correction in a finite trap geometry and induce the anomalous enhancement of the damping rate.
@@
Figure 1(b) shows results for the smallest finite relaxation rates $\tau_{m}$ with odd and even parity, where we use the expansion method developed in Refs. [20, 21] to determine the eigenvalues of the collision integral to arbitrary precision. The key observation is the parity-dependence of the lifetime at large $\beta\mu$, where Pauli blocking leads to a different scaling of the odd- and even-lifetimes with chemical potential and temperature: odd modes scale as $\beta^{4}\mu^{3}$, while even modes scale as $\beta^{2}\mu$ [indicated in Fig. 1(b)]. Current experiments reach $T/T_{F}\approx 0.1-0.5$ [14, 35, 36, 37, 38], which is indicated by the dashed grey line in Fig. 1(b) for the BCS limit, for which the chemical potential is approximately that of a non-interacting Fermi gas. Already at this point the odd and even lifetimes differ by a factor of two, and even a minor further decrease in temperature (i.e., larger $\beta\mu$) will significantly enhance the odd-even effect: For example, for $T/T_{F}\approx 0.05$ the odd and even lifetimes differ by one order of magnitude. On the BEC side [purple shaded area in Fig. 1(b)], the chemical potential becomes negative, $\beta\mu<0$, and scales with the binding energy $\epsilon_{b}=1/(m^{*}a_{2}^{2})$ of the two-body bound state, $\mu=-\epsilon_{b}/2$. In this regime, odd and even relaxation times are approximately equal and scale as $\tau_{m}\sim e^{-\beta\mu}$. As discussed, intuitively on the BEC side there is no well-defined Fermi surface and thus one expects the odd-even effect to vanish. This result implies that the odd-even effect is tuned not only by changing the temperature but also the interaction strength. Since the odd-even effect is due to the kinematic constraints of low-energy scattering in the presence of a Fermi surface, it is not overly sensitive to the structure of the matrix element [18, 20]. We have confirmed this by considering various approximations for the two-body scattering matrix element in the collision integral, see the supplemental material [39] for details.
@@
Shear viscosity We now propose to observe the odd-even effect along the BCS-BEC crossover by examining the transverse collective dynamics of the gas in real space. In atomic gases, a transverse flow is induced by collective quadrupole modes, which are anisotropic oscillations of a harmonically trapped quantum gas [40, 41, 42, 43, 44, 45, 14, 46, 35, 47, 48]. This is shown schematically in Fig. 1(c) alongside the radial equilibrium density distribution for the Fermi gas for different trap confinements, which at low temperatures is a Thomas-Fermi profile with radius $R=\hbar k_{F}/(m^{*}\omega_{r})=\sqrt{2\hbar\sqrt{N}/(m^{*}\omega_{r})}$. The quadrupolar flow is incompressible and there is no change in the internal energy [40, 41, 42, 14]. This yields an oscillation frequency that is independent of the equation of state in the strongly interacting limit, $\omega_{Q}=\sqrt{2}\omega_{r}$ [14, 49], where $\omega_{r}$ is the frequency of the harmonic trap.
@@
The damping rate $\Gamma_{Q}$ of the quadrupole mode is linked to the shear viscosity spectral function evaluated at the quadrupole mode frequency, $\eta_{\bf r}(\omega_{Q},{\bf q})$, via [43, 44, 45, 14, 46, 35]
@@
In the limit of very shallow harmonic traps, the system is nearly homogeneous and the damping rate is related to the shear viscosity spectral function at $q=0$. This is the hydrodynamic limit that is dominated by the relaxation of the even-parity $m=\pm 2$ mode [32, 33, 34]. At finite trap frequencies, i.e., for finite trap curvature $q\sim 1/R$, the odd-parity modes ($m=\pm 3$ and higher angular momenta) contribute and there is a substantial difference between the shear viscosity spectral function $\eta(\omega_{Q},q)$, and its hydrodynamic counterpart $\eta_{\rm HD}=\eta(\omega_{Q},q=0)$. For fixed particle number, the relevant wavenumber $q\sim\sqrt{\omega_{r}}/N^{1/4}$ and frequency $\omega_{Q}=\sqrt{2}\omega_{r}$ only depend on the harmonic trap frequency, $\omega_{r}$. Hence, by tuning the trap frequency, we explore the shear viscosity spectral function on a line in the $\{\omega,q\}$ parameter space. In Fig. 1(d), we compare the results for the shear viscosity at finite and zero momentum for $T/T_{F}=0.1$ following the Fermi liquid approach outlined in the methods section and the supplemental material [39]. Already for intermediate trap sizes, we see a noticeable difference between the hydrodynamic and finite momentum predictions, which indicates a significant enhancement of the damping rate compared to the hydrodynamic prediction that is due to the anomalous increase in the lifetime of odd-parity modes.

### v2

@@ Abstract
Two-dimensional electron gases are predicted to possess an anomalous “tomographic” transport regime that is marked by an odd-even effect in the relaxation times, with odd-parity deformations of the Fermi surface becoming long-lived in comparison to even-parity ones. In this work, we establish that neutral two-component atomic Fermi gases also exhibit this tomographic effect. By diagonalizing the Fermi liquid collision integral, we identify odd-parity modes with anomalously long lifetimes below temperatures $T\leq 0.15T_{F}$, which is within reach of cold atom experiments. Furthermore, in contrast to electron gases, we find that the odd-even effect in neutral gases is widely tuneable with interactions along the BCS-BEC crossover and is suppressed on the BEC side. We propose as an experimental signature of the odd-even effect the damping rate of quadrupole oscillations, which is anomalously enhanced due to the presence of long-lived odd-parity modes. Our findings suggest that the dynamics of two-dimensional Fermi gases is richer than previously thought and should include additional long-lived modes.
@@ Introduction
Thermodynamic and transport properties of the normal state of metals as well as 3He are well described by a semiclassical Fermi gas model [1, 2]. The microscopic justification of this surprisingly simple description, despite nominally strong interactions, is provided by Landau’s theory of the Fermi liquid [3, 4], which describes thermodynamic and transport properties in terms of a collection of weakly interacting quasiparticles [5, 6, 7]. The validity of this picture relies on the presence of a Fermi surface and concomitant Pauli blocking, which strongly restricts the available phase space for quasiparticle scattering, leading to well-defined quasiparticles with lifetimes that diverge quadratically at low temperatures, $\tau\sim 1/T^{2}$. In recent years, Fermi liquid theory has also been applied to a variety of ultracold quantum gas setups, such as minority polarons in strongly spin-imbalanced Fermi gases [8, 9, 10, 11], the normal phase of balanced Fermi gases above the superfluid transition [12, 13, 14, 15], transport and collective modes [16] and even few-body systems [17]. For transport measurements, quasiparticle interactions lead to internal dissipative processes with relaxation times that set the magnitude of transport coefficients like the shear viscosity. The general expectation is that these transport times are equal to the quasiparticle lifetimes.
@@
Figure 1: Tomographic transport in atomic gases. (a) Nonequilibrium momentum-space distribution for different temperatures $T$ (where $\beta=1/[k_{B}T]$ is the inverse temperature and $k_{B}$ is the Boltzmann constant) and (b) dimensionless relaxation rates $\beta\hbar/\tau_{m}$ of the first collective even-parity Fermi surface deformation with angular momentum $m=2$ (blue) and the first collective odd-parity deformation with $m=3$ (orange). Even (odd) parity refers to an equal (opposite) deformation of the distribution at opposite momenta. With decreasing temperature and increasing density (i.e., increasing $\beta\mu$, where $\mu$ is the chemical potential), the equilibrium Fermi surface becomes sharper and Pauli blocking diminishes the available phase space for scattering, increasing the lifetime of the modes. The conventional Fermi liquid relaxation rate $\sim T^{2}/\mu$ of the even-parity mode is set by an angular redistribution on the Fermi surface of pairs with equal and opposite momenta. Such a process does not relax the odd-parity modes, however, which have a significantly slower relaxation rate $\sim T^{4}/\mu^{3}$. For interacting Fermi gases along the BCS-BEC crossover, the discrepancy of odd and even-parity lifetimes is present on the BCS side of the crossover, for which the chemical potential is positive, $\beta\mu>0$ (the gray dashed line indicates the non-interacting chemical potential at a moderate small temperature $T/T_{F}=0.1$ for reference). The effect is then reduced with increasing pairing interaction and vanishes on the BEC side where $\beta\mu<0$. (c) Real-space dynamics and (d) damping rate $\Gamma_{Q}$ as a function of trap frequency $\omega_{r}$ of the quadrupole mode in a harmonic trap. The quadrupole mode oscillates with frequency $\omega_{Q}=\sqrt{2}\omega_{r}$ and involves a shear flow in real space. The real-space density profile is shown in the insets and sets the equilibrium size $R$. The damping rate of the quadrupole mode is controlled by the shear viscosity spectral function $\eta(\omega,q)$, which is proportional to the relaxation time of a collective shear deformation of the Fermi surface in momentum space. In a bulk system (shallow trap), this deformation corresponds to the hydrodynamic even-parity mode with $m=2$, but at finite confinement odd-parity modes contribute, and the increased odd-parity lifetime leads to an increased damping rate (orange line) compared to the hydrodynamic prediction (blue line). This damping provides evidence of the enhanced odd-parity lifetime.
@@
Recent theoretical work in the context of interaction-dominated electron gases in two dimensions has provided evidence to the contrary and indicates the existence of deformations of the Fermi surface that have significantly longer lifetimes [18, 19, 20, 21, 22]: Indeed, a standard Fermi liquid scaling $\tau_{\rm e}\sim 1/T^{2}$ is argued to only apply to a subclass of collective quasiparticle deformations—dubbed even-parity modes—while the remaining modes—dubbed odd-parity modes—decay significantly more slowly as $\tau_{\rm o}\sim 1/T^{4}$. This leads to the exciting perspective that, at low temperatures, these long-lived modes impact the hydrodynamic description of the gas (in addition to modes that are linked to conserved currents and have infinite lifetime), such that the transport properties of Fermi liquids become richer than commonly assumed. For this reason, there is now intense effort to identify experimental signatures of odd-parity transport, or “tomographic transport”, in electron Fermi liquids [19, 23, 24, 25, 26, 27]. Quite generally, an isolated odd-parity response that is not dominated by even-parity modes appears in transverse probes at finite wavelengths, such as the transverse conductivity [19] or transverse collective modes [23]. However, it has so far not been possible to detect such transverse probes in electron Fermi liquids, while in 3He, measurements of transverse sound have been inconclusive [28, 29]. By contrast, quantum gases offer an entirely separate way of inducing transverse dynamics by manipulating the real-space confinement of an external trap. Moreover, since on a microscopic level the odd-even effect is linked to the Fermi statistics and Pauli blocking, one should expect (and as we confirm theoretically) that this effect is not only present for electron gases but also in charge-neutral Fermi liquids. The purpose of this work is thus to initiate a search for the odd-even effect in neutral quantum gases.
@@
The key results of our work, illustrated in Fig. 1, are twofold: First, we establish that, just as for electron gases, an enhancement in the lifetime of odd-parity deformations of the Fermi surface also exists for neutral atomic Fermi quantum gases [Fig. 1(a)]. From an exact diagonalization of the linearized kinetic collision integral (see methods), we find that these lifetimes differ significantly within temperature ranges accessible in current cold atom experiments [Fig. 1(b)]. Moreover, in a marked difference from electron gases, we demonstrate that the effect is widely tunable with interaction strength along the BCS-BEC crossover and vanishes near resonance and on the BEC side of the crossover where the chemical potential $\mu$ is negative and the Fermi surface disappears. As a second key result, we propose as an experimental signature of the odd-even effect the damping rate of collective quadrupole oscillations in a harmonic trap, which involve a shear flow of atoms in real space [Fig. 1(c)]. The damping rate of the oscillations is governed by the shear viscosity, which is evaluated at a finite wavenumber set by the trap curvature. As discussed above, this implies that the damping rate is proportional to the relaxation time of the momentum space deformation that is induced by the shear flow in the harmonic trap, $\eta\approx np_{F}v_{F}\tau_{\eta}$. For a shallow trap, the standard hydrodynamic result for the shear viscosity involves an even-parity deformation with conventional Fermi liquid lifetime, while for larger confinements, odd-parity modes dominate. Since the shear viscosity is related to the lifetime of a deformation of the Fermi surface induced by the shear flow, the reduced scattering of the odd-parity modes translates to an anomalously enhanced shear viscosity at finite wavenumber and hence an increased damping rate of the quadrupole mode when odd-parity modes become relevant. Indeed, a full evaluation of the momentum- and frequency-dependent shear viscosity spectral function [Fig. 1(d)] shows a strong increase of the damping rate compared to the hydrodynamic prediction for intermediate trap frequencies $\omega_{r}$ and realistic system parameters. Moreover, as will be discussed below, the effect is enhanced even further both at low temperatures and for small atom numbers, i.e., for mesoscopic Fermi gases. Overall, these findings imply that in experiments with cold gases, the odd parity effect can be widely tuned in a characteristic manner with interaction strength, temperature, harmonic trap frequency, and particle number. Ultracold Fermi gases are thus an ideal platform to study the anomalous relaxation dynamics of collective modes, and it should be readily possible to establish the odd-even effect in the quasiparticle lifetimes by comparing the damping of the quadrupole mode at different experimental parameter values.
@@ Results
Quasiparticle lifetimes We first establish the odd-even effect in the lifetimes of the odd and even parity modes of the quasiparticle distribution function at low temperatures on the BCS side of the normal region of the BCS-BEC crossover (where a well-defined Fermi surface exists) and show that it vanishes along the crossover. Intuitively, this is linked to the fact that fermions along the crossover form increasingly tightly bound dimers, which reduces the chemical potential and destroys the Fermi surface, and with it the Pauli blocking that is at the heart of the odd-even effect. For 2D atomic gases, interactions are short-ranged and described by a delta-function with strength $\hbar^{2}\tilde{g}/m^{*}$ with a dimensionless parameter $\tilde{g}=-2\pi/\ln(k_{F}a_{2})$, where $k_{F}$ is the Fermi wavevector, $m^{*}$ is the mass of the atoms and $a_{2}$ is the 2D scattering length. At a given temperature and density, the interactions interpolate between the BCS side $\tilde{g}<0$ or $\ln(k_{F}a_{2})>0$] and the BEC side $\tilde{g}>0$ or $\ln(k_{F}a_{2})<0$], tuned by either Feshbach  [30] or confinement-induced resonances [31].
Figure 2: Shear viscosity spectral function and anomalous damping of the quadrupole mode. (a) Shear viscosity spectral function in units of $\hbar n$, where $n$ is the particle density at the center of the harmonic trap, as a function of wave number $q$ and frequency $\omega$ (both in thermal units with $\lambda_{T}=\sqrt{2\pi\beta\hbar^{2}/m^{*}}$). The white solid (dashed) line describes the path in parameter space ($q=1/R,\omega=\omega_{Q}$) explored by tuning the trap confinement for $N_{1}=10^{3}$ ($N_{2}=10^{6}$) atoms. (b)-(d): Ratio of the shear viscosity spectral function $\eta(\omega=\omega_{Q},q=1/R)$ and the hydrodynamic shear viscosity $\eta_{HD}=\eta(\omega_{Q},0)$ as a function of (b) trap frequency, (c) interaction strength, and (d) particle number. Unless otherwise specified, we use an atom number $N=10^{3}$ and interaction strength $\ln(k_{F}a_{2})=0.63$ (corresponding to $\beta\mu=9$). In each case, we observe a pronounced deviation from the hydrodynamic prediction for smaller atom number, interactions that are deeper in the BCS limit, and for weaker harmonic traps (or lower temperatures), respectively.
The fundamental quantities to describe are the relaxation rates of collective deviations in the quasiparticle distribution from local equilibrium (which is given by a Fermi-Dirac distribution $n_{\rm FD}({\bf p})$). These relaxation rates are defined as the eigenvalues of the linearized collision integral (see methods),
@@
Due to the rotational invariance of the system, the eigenmodes $\psi_{m}$ carry a definite angular momentum quantum number $m=0,\pm 1,\pm 2$.... Modes with even (odd) $m$ have even (odd) parity [see Fig. 1(a)]. An additional radial quantum number (not indexed explicitly here) describes the number of radial nodes in the eigenfunction, with faster relaxation (i.e., larger $1/\tau_{m}$) for modes with increasing node number [21]. Due to conservation of particle number, energy, and momentum, the lowest eigenvalues with $m=0$ and $m=\pm 1$ are zero. The leading finite even-parity mode has then angular momentum $m=\pm 2$ while the leading odd-parity mode has angular momentum $m=\pm 3$ [21]. The corresponding momentum space deformations are shown in Fig. 1(a) for three different temperatures. The $m=2$ mode sets the dominant contribution to the hydrodynamic shear viscosity [32, 33, 34], while the second mode will give the leading correction in a finite trap geometry and induce the anomalous enhancement of the damping rate.
@@
Figure 1(b) shows results for the smallest finite relaxation rates $\tau_{m}$ with odd and even parity, where we use the expansion method developed in Refs. [20, 21] to determine the eigenvalues of the collision integral to arbitrary precision. The key observation is the parity-dependence of the lifetime at large $\beta\mu$ ($\beta=1/[k_{B}T]$ and $k_{B}$ is Boltzmann’s constant), where Pauli blocking leads to a different scaling of the odd- and even-lifetimes with chemical potential and temperature: odd modes scale as $\beta^{4}\mu^{3}$, while even modes scale as $\beta^{2}\mu$ [indicated in Fig. 1(b)]. Current experiments reach $T/T_{F}\approx 0.1-0.5$ [14, 35, 36, 37, 38], which is indicated by the dashed gray line in Fig. 1(b) for the BCS limit, for which the chemical potential is approximately that of a non-interacting Fermi gas. Already at this point the odd and even lifetimes differ by a factor of two, and even a minor further decrease in temperature (i.e., larger $\beta\mu$) will significantly enhance the odd-even effect: For example, for $T/T_{F}\approx 0.05$ the odd and even lifetimes differ by one order of magnitude. On the BEC side [purple shaded area in Fig. 1(b)], the chemical potential becomes negative, $\beta\mu<0$, and scales with the binding energy $\epsilon_{b}=1/(m^{*}a_{2}^{2})$ of the two-body bound state, $\mu=-\epsilon_{b}/2$. In this regime, odd and even relaxation times are approximately equal and scale as $\tau_{m}\sim e^{-\beta\mu}$. As discussed, intuitively on the BEC side there is no well-defined Fermi surface and thus one expects the odd-even effect to vanish. This result implies that the odd-even effect is tuned not only by changing the temperature but also the interaction strength. Since the odd-even effect is due to the kinematic constraints of low-energy scattering in the presence of a Fermi surface, it is not overly sensitive to the structure of the two-particle matrix element in the collision integral [18, 20]. We have confirmed this by considering various approximations for the two-body scattering matrix, see Supplementary Note I for details.
@@
Shear viscosity We now propose to observe the odd-even effect along the BCS-BEC crossover by examining the transverse collective dynamics of the gas in real space. In atomic gases, a transverse flow is induced by collective quadrupole modes, which are anisotropic oscillations of a harmonically trapped quantum gas [39, 40, 41, 42, 43, 44, 14, 45, 35, 46, 47]. This is shown schematically in Fig. 1(c) alongside the radial equilibrium density distribution for the Fermi gas for different trap confinements, which at low temperatures is a Thomas-Fermi profile with radius $R=\hbar k_{F}/(m^{*}\omega_{r})=\sqrt{2\hbar\sqrt{N}/(m^{*}\omega_{r})}$. The quadrupolar flow is incompressible and there is no change in the internal energy [39, 40, 41, 14]. This yields an oscillation frequency that is independent of the equation of state in the strongly interacting limit, $\omega_{Q}=\sqrt{2}\omega_{r}$ [14], where $\omega_{r}$ is the frequency of the harmonic trap. In general, the frequency of the quadrupole mode depends on the interaction strength and scales as $\omega_{Q}=\sqrt{2(2+\tilde{g})/(1-\tilde{g})}\omega_{r}$ in the collisionless limit [39]. However, this change in the frequency does not qualitatively affect our analysis, which relies primarily on the finite wavenumber $q$.
@@
The damping rate $\Gamma_{Q}$ of the quadrupole mode is linked to the local shear viscosity spectral function evaluated at the quadrupole mode frequency, $\eta(\omega_{Q},{\bf r})$, via [42, 43, 44, 14, 45, 35]
$\Gamma_{Q}$ $=E_{{\rm kin},Q}^{-1}\int d{\bf r}\,\eta(\omega_{Q},{\bf r})\frac{1}{2}\Bigl{(}\partial_{x}{\bf v}_{Q}^{y}+\partial_{y}{\bf v}_{Q}^{x}\Bigr{)}^{2}.$ (3)
@@
In the limit of very shallow harmonic traps, the system is nearly homogeneous and the damping rate is related to the shear viscosity spectral function at $q=0$. This is the hydrodynamic limit that is dominated by the relaxation of the even-parity $m=\pm 2$ mode [32, 33, 34]. At finite trap frequencies, i.e., for finite trap curvature $q\sim 1/R$, the odd-parity modes ($m=\pm 3$ and higher angular momenta) contribute. To obtain an estimate of the characteristic parameter scales at which these odd-parity modes become relevant, we use a relaxation time approximation with a single characteristic even (odd) parity modes decay rate $\gamma_{e}$ ($\gamma_{o}$). At low temperatures, these relaxation rates scale as $\gamma_{e}\sim\tilde{g}^{2}T^{2}/T_{F}$ and $\gamma_{o}\sim\tilde{g}^{2}T^{4}/T_{F}^{3}$ [19, 20, 21]. When the odd-even effect is present $\gamma_{e}\gg\gamma_{o}$, there is an emergent length scale that defines the tomographic transport limit $\xi=v_{F}/\sqrt{\gamma_{e}\gamma_{o}}$, where for $q\xi\sim 1$ the system is in the tomographic transport regime, while for $q\xi\ll 1$ is the hydrodynamic regime. In the presence of a harmonic trap, both the relevant wavenumber $q\sim\sqrt{\omega_{r}}/N^{1/4}$ and frequency $\omega_{Q}=\sqrt{2}\omega_{r}$ are set by the harmonic trap frequency $\omega_{r}$ and the atom number $N$. Hence, by tuning the trap frequency we explore the shear viscosity spectral function on a line in the $\{\omega,q\}$ parameter space. Expressed in terms of the trap frequency, the condition $q\xi\sim 1$ to observe the odd-even effect in a harmonic trap translates to
$q\xi\sim\frac{N\omega_{r}^{3}}{\tilde{g}^{2}T^{3}}\sim 1,$ (4)
implying that the odd-even effect will be important for $\omega_{r}\sim\tilde{g}^{2}T^{3}/N$.
In Fig. 1(d), we compare the results for the shear viscosity at finite and zero momentum for $T/T_{F}=0.1$ following the Fermi liquid approach outlined in the methods section and Supplementary Note II. Already for intermediate trap sizes, we see a noticeable difference between the hydrodynamic and finite momentum predictions, which indicates a significant enhancement of the damping rate compared to the hydrodynamic prediction that is due to the anomalous increase in the lifetime of odd-parity modes.